At a gathering consisting of 10 men and 20 women, two door prizes are awarded. Find the probability that both prizes are won by men. The winning ticket is not replaced
Answer by rothauserc(4718)

There are 10 + 20 = 30 people
:
Probability both tickets are won by men = (10/30) * (9/29) = 3/29
: